OpenStudy (anonymous):

find the midpoint of the line segment joining the points p1 and p2 p1=(-5,4) p2=(1,6)

OpenStudy (mertsj):

(-5+1)/2=???

OpenStudy (anonymous):

-2

OpenStudy (mertsj):

\[\frac{-5+1}{2}=\frac{-4}{2}=-2\]

OpenStudy (mertsj):

yep Now do the y's

OpenStudy (anonymous):

5

OpenStudy (mertsj):

\[\frac{4+6}{2}=\frac{10}{2}=\]

OpenStudy (mertsj):

yes. so the midpoint is (-2,5)

OpenStudy (mertsj):

Let's see if that makes any sense.

OpenStudy (anonymous):

That wasn't hard! Thanks for showing me!

OpenStudy (mertsj):